A valószínűségszámítás gyakorlaton való részvétel

dátum:  II. 17. II. 24. III. 2. III. 9. III. 16. III. 23. III. 30.

diákok:

Angyal Orsolya  +  +  +  +  +
Balog Erzsébet  +  +  −  +  −
Biró Ildikó  +  +  −  −  −
Bujdosó Tünde  +  −  +  −  +
Erdős Veronika  +  −  −  −  +
Juhász Mónika  +  −  +  +  −
Molnár Anita  +  +  −  +
Nagy Gábor Bertalan  +  +  −  −  −
Palik Erika  +  +  +  +  −
Popovics Angéla  +  +  +  +  +
Radics Ágnes  +  +  +  −  +
Szánkai Nikolett  +  −  +  −  −
Templom Mária  +  +  +  +  −
Vattamány Magdolna  +  +  +  −  −